# Temperature Dependence of Heavy Quark Diffusion from (2+1)-flavor Lattice QCD

### v1

@@ Appendix C Quasi-particle model for charm and bottom agrees of freedom
At sufficiently high temperatures, the partial pressures of charm mesons and baryons are expected to vanish, i.e., $p_{C}^{M}\simeq p_{C}^{B}\simeq 0$, while the charm-quark contribution $p_{C}^{q}$ approaches its ideal gas limit. Lattice QCD results confirm that $p_{C}^{M}(T)$ and $p_{C}^{B}(T)$ become small for $T>190$ MeV, and the total charm pressure is then dominated by the quark contribution [73, 74, 75]. Thus, at these temperatures, the generalized charm susceptibilities can be accurately described by a charm-quark gas.
@@
In this study, we use continuum extrapolated lattice QCD results for the second and fourth order charm susceptibilities as calculated in Refs. [76] and [75]. These susceptibilities are defined as:
$\chi_{n}^{C}=\frac{\partial^{n}(p_{C}(T)/T^{4})}{\partial(\mu_{C}/T)^{n}},~{}n=2,4\,.$ (C.2)
@@
Here $\mu_{C}$ is the charm chemical potential. Due to the large charm-quark mass, we can use Boltzmann statistics, for which $\chi_{2}^{C}\equiv\chi_{c}=\chi_{4}^{C}=p_{C}(T)$. This relation is supported by lattice QCD results [72].
@@
We use interpolations to obtain values of the susceptibilities at the temperatures relevant to our analysis. Figure 24 shows these interpolated values alongside the predictions from the QM-HRG. Specifically in the figure, we highlight the interpolated value of the charm susceptibility at $T=188$ MeV used in our study. While the QM-HRG model does not fully describe the lattice results for the generalized susceptibilities across all temperatures [75], it is noteworthy that the susceptibilities agree within errors with the QM-HRG predictions up to approximately $188$ MeV.
@@
Since charm quarks dominate the relevant degrees of freedom for $T>190$ MeV, we use a simple quasi-particle model defined by Eqs. (5.8), (5.9), and (5.10). In this regime, the quasi-particle mass $M_{c}(T)$ is determined directly from the lattice QCD results for the charm susceptibility $\chi_{2}^{C}$, as shown in Fig. 24 and listed in Tab. 3. Using these values of $M_{c}(T)$, we then compute the mean squared thermal velocity ${\langle\mathbf{v}^{2}\rangle}$ MeV and mean squared thermal momentum ${\langle\mathbf{p}^{2}\rangle}$ for temperatures at and above $195$ MeV. The corresponding uncertainties are determined by the statistical errors on $\chi_{c}$. This is the same approach used in the previous HotQCD study [21].
@@
For temperatures below $190$ MeV, we instead use the lattice QCD results for the charm-quark partial pressure $p_{C}^{q}(T)$ from Ref. [75] to determine the quasi-particle mass. This is done by using Eq. (5.8) and replacing $\chi_{c}$ with $p_{C}^{q}$. In this low-temperature regime, charm meson and charm baryon contributions become relevant and must be included. For this purpose, we adopt the 1S1P-HRG model introduced in Ref. [75], which includes the ground-state charm hadrons and their parity partners (first orbital excitations) [84, 85]. This model has been shown to give a fair description of the lattice results for $p_{C}^{M}$ and $p_{C}^{B}$ [75]. Using the charm-quark mass $M_{c}(T)$ extracted from $p_{C}^{q}(T)$ and the hadron masses listed in the 1S1P-HRG model, we calculate ${\langle\mathbf{v}^{2}\rangle}$ and ${\langle\mathbf{p}^{2}\rangle}$ for $T<190$ MeV. The results are summarized in Tab. 3 and are referred to as Model 1, with corresponding quantities labeled using the subscript $c_{1}$. For sufficiently large masses, the relation ${\langle\mathbf{p}^{2}\rangle}=3M_{c}T$ holds. Therefore, in the table we report the scaled quantity $\hat{\langle\mathbf{p}^{2}\rangle}={\langle\mathbf{p}^{2}\rangle}/(3M_{c}T)$ to show the proximity of our results with heavy-mass approximation. In fact, in order to determine $D_{s}$ from $\kappa$ we only need the value of ${\langle\mathbf{p}^{2}\rangle}/(3M_{c}T)$, cf. Eq. (2.9).
@@
To compute $\chi_{b}$ within QM-HRG, we include the known open-beauty hadrons listed in the Particle Data Group, and supplement the spectrum with additional states predicted by relativistic quark model calculations [84, 85], as done for the charm sector. Above the crossover temperature, we assume the same set of bottom degrees of freedom as for charm, and write the total bottom susceptibility as:
$\chi_{b}(T)=p_{B}(T)=p_{B}^{M}(T)+p_{B}^{B}(T)+p_{B}^{q}(T)\,,$ (C.3)
where $p_{B}^{M}(T)$, $p_{B}^{B}(T)$ and $p_{B}^{q}(T)$ represent the partial pressures from bottom quarks, mesons, and baryons, respectively. For the hadronic contributions $p_{B}^{M}(T)$ and $p_{B}^{B}(T)$, we use the 1S1P-HRG model, which again includes ground-state open-beauty hadrons and their first orbital excitations. For the spectrum of beauty hadrons we use the masses from Particle Data Group and relativistic quark model [84, 85]. The corresponding pressures in QM-HRG and 1S1P-HRG are shown in Fig. 25. With $\chi_{b}(T)$ from QM-HRG and $p_{B}^{M}(T)$ and $p_{B}^{B}(T)$ from 1S1P-HRG, we isolate the quark contribution $p_{B}^{q}(T)$, and use Eq. (5.8) replacing $\chi_{b}(T)\rightarrow p_{B}^{q}(T)$ on the left hand side to extract the quasi-particle bottom-quark mass $M_{b}(T)$. The resulting values are listed in Table  4.
@@
We note that the mean squared thermal velocity for charm and bottom quarks can also be estimated using the NLO results for the massive vector correlation function [86], as discussed in Ref. [87], when the temperature is sufficiently large. For the temperature range $T=251-352$ MeV, we find that the resulting estimates for ${\langle\mathbf{v}^{2}\rangle}_{c}$ and ${\langle\mathbf{v}^{2}\rangle}_{b}$ are lower by approximately 5–10% and 15–17%, respectively, compared to our quasi-particle model results.
In Fig. 27, we summarize the temperature dependence of ${\langle\mathbf{v}^{2}\rangle}$ and ${\langle\mathbf{p}^{2}\rangle}$ for charm and bottom degrees of freedom, along with their estimated uncertainties, as a function of $T/M$. The shades on the data represent the systematic differences that arise from comparing Models 2 and 3 to Model 1. At leading order in the $T/M$ expansion, we have the simple relations ${\langle\mathbf{v}^{2}\rangle}=3T/M$ and ${\langle\mathbf{p}^{2}\rangle}=3MT$. Our results show significant deviations from these LO predictions. However, including the NLO terms in the expansion in $T/M$ [27] substantially improves the agreement, as can be seen in the figure.

### v2

@@ Appendix C Quasi-particle model for charm and bottom degrees of freedom
At sufficiently high temperatures, the partial pressures of charm mesons and baryons are expected to vanish, i.e., $p_{C}^{M}\simeq p_{C}^{B}\simeq 0$, while the charm-quark contribution $p_{C}^{q}$ approaches its ideal gas limit. Lattice QCD results confirm that $p_{C}^{M}(T)$ and $p_{C}^{B}(T)$ become small for $T>190$ MeV, and the total charm pressure is then dominated by the quark contribution Mukherjee:2015mxc ; Bazavov:2023xzm ; Kaczmarek:2025dqt . Thus, at these temperatures, the generalized charm susceptibilities can be accurately described by a charm-quark gas.
@@
In this study, we use continuum extrapolated lattice QCD results for the second and fourth order dimensionless charm susceptibilities as calculated in Refs. Bellwied:2015lba and Kaczmarek:2025dqt , respectively. These susceptibilities are defined as:
$\hat{\chi}_{n}^{C}=\frac{\partial^{n}(p_{C}(T)/T^{4})}{\partial(\mu_{C}/T)^{n}},~n=2,4\,.$ (C.2)
@@
Here $\mu_{C}$ is the charm chemical potential. Due to the large charm-quark mass, we can use Boltzmann statistics, for which $\hat{\chi}_{4}^{C}=\hat{\chi}_{2}^{C}\equiv\chi_{c}/T^{2}=p_{C}(T)/T^{4}$. This relation is supported by lattice QCD results Bazavov:2014yba .
@@
We use interpolations to obtain values of the susceptibilities at the temperatures relevant to our analysis. Figure 24 shows these interpolated values alongside the predictions from the QM-HRG. Specifically in the figure, we highlight the interpolated value of the charm susceptibility at $T=188$ MeV used in our study. While the QM-HRG model does not fully describe the lattice results for the generalized susceptibilities across all temperatures Kaczmarek:2025dqt , it is noteworthy that the susceptibilities agree within errors with the QM-HRG predictions up to approximately $188$ MeV.
@@
Since charm quarks dominate the relevant degrees of freedom for $T>190$ MeV, we use a simple quasi-particle model defined by Eqs. (5.9), (5.10), and (5.11). In this regime, the quasi-particle mass $M_{c}(T)$ is determined directly from the lattice QCD results for the charm susceptibility $\hat{\chi}_{2}^{C}$, as shown in Fig. 24 and listed in Tab. 3. Using these values of $M_{c}(T)$, we then compute the mean squared thermal velocity $\langle\mathbf{v}^{2}\rangle$} and mean squared thermal momentum $\langle\mathbf{p}^{2}\rangle$} for temperatures at and above $195$ MeV. The corresponding uncertainties are determined by the statistical errors on $\hat{\chi}_{2}^{C}$. This is the same approach used in the previous HotQCD study Altenkort:2023eav .
@@
For temperatures below $190$ MeV, we instead use the lattice QCD results for the charm-quark partial pressure $p_{C}^{q}(T)$ from Ref. Kaczmarek:2025dqt to determine the quasi-particle mass. This is done by using Eq. (5.9) and replacing $\chi_{c}$ with $p_{C}^{q}/T^{2}$. In this low-temperature regime, charm meson and charm baryon contributions become relevant and must be included. For this purpose, we adopt the 1S1P-HRG model introduced in Ref. Kaczmarek:2025dqt , which includes the ground-state charm hadrons and their parity partners (first orbital excitations) Ebert:2011kk ; Chen:2022asf . This model has been shown to give a fair description of the lattice results for $p_{C}^{M}$ and $p_{C}^{B}$ Kaczmarek:2025dqt . Using the charm-quark mass $M_{c}(T)$ extracted from $p_{C}^{q}(T)$ and the hadron masses listed in the 1S1P-HRG model, we calculate $\langle\mathbf{v}^{2}\rangle$} and $\langle\mathbf{p}^{2}\rangle$} for $T<190$ MeV. The results are summarized in Tab. 3 and are referred to as Model 1, with corresponding quantities labeled using the subscript $c_{1}$. For sufficiently large masses, the relation ${\langle\mathbf{p}^{2}\rangle}=3M_{c}T$ holds. Therefore, in the table we report the scaled quantity $\hat{\langle\mathbf{p}^{2}\rangle}={\langle\mathbf{p}^{2}\rangle}/(3M_{c}T)$ to show the proximity of our results to the heavy-mass approximation. In fact, in order to determine $D_{s}$ from $\kappa$ we only need the value of ${\langle\mathbf{p}^{2}\rangle}/(3M_{c}T)$, cf. Eq. (2.9).
@@
To compute $\chi_{b}$ within QM-HRG, we include the known open-beauty hadrons listed in the Particle Data Group, and supplement the spectrum with additional states predicted by relativistic quark model calculations Ebert:2011kk ; Chen:2022asf , as done for the charm sector. Above the crossover temperature, we assume the same set of bottom degrees of freedom as for charm, and write the total bottom susceptibility as:
$\chi_{b}(T)\cdot T^{2}=p_{B}(T)=p_{B}^{M}(T)+p_{B}^{B}(T)+p_{B}^{q}(T)\,$, (C.3)
where $p_{B}^{M}(T)$, $p_{B}^{B}(T)$ and $p_{B}^{q}(T)$ represent the partial pressures from bottom mesons, bottom baryons, and bottom quarks respectively. For the hadronic contributions $p_{B}^{M}(T)$ and $p_{B}^{B}(T)$, we use the 1S1P-HRG model, which again includes ground-state open-beauty hadrons and their first orbital excitations. For the spectrum of beauty hadrons we use the masses from Particle Data Group and relativistic quark model Ebert:2011kk ; Chen:2022asf . The corresponding pressures in QM-HRG and 1S1P-HRG are shown in Fig. 25. With $\chi_{b}(T)$ from QM-HRG and $p_{B}^{M}(T)$ and $p_{B}^{B}(T)$ from 1S1P-HRG, we isolate the quark contribution $p_{B}^{q}(T)$, and use Eq. (5.9) replacing $\chi_{b}(T)\rightarrow p_{B}^{q}(T)$ on the left hand side to extract the quasi-particle bottom-quark mass $M_{b}(T)$. The resulting values are listed in Table  4.
@@
We note that the mean squared thermal velocity for charm and bottom quarks can also be estimated using the NLO results for the massive vector correlation function Burnier:2012ze , as discussed in Ref. Laine_massive_VC , when the temperature is sufficiently large. For the temperature range T=251352T=251-352 MeV, we find that the resulting estimates for ${\langle\mathbf{v}^{2}\rangle}_{c}$ and ${\langle\mathbf{v}^{2}\rangle}_{b}$ are lower by approximately 5–10% and 15–17%, respectively, compared to our quasi-particle model results.
In Fig. 27, we summarize the temperature dependence of $\langle\mathbf{v}^{2}\rangle$} and $\langle\mathbf{p}^{2}\rangle$} for charm and bottom degrees of freedom, along with their estimated uncertainties, as a function of $T/M$. The shades on the data represent the systematic differences that arise from comparing Models 2 and 3 to Model 1. At leading order in the $T/M$ expansion, we have the simple relations ${\langle\mathbf{v}^{2}\rangle}=3T/M$ and ${\langle\mathbf{p}^{2}\rangle}=3MT$. Our results show significant deviations from these LO predictions. However, including the NLO terms in the expansion in $T/M$ Bouttefeux:2020ycy substantially improves the agreement, as can be seen in the figure.
@@
Acknowledgements.
This material is based on work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Contract No. DE-SC0012704 and by the Scientific Discovery through Advanced Computing (SciDAC) award "Fundamental Nuclear Physics at the Exascale and Beyond" and the Topical Collaboration in Nuclear Theory "Heavy-Flavor Theory (HEFTY) for QCD Matter". Part of the analysis contained in this publication was based on the AnalysisToolBox code AnalysisToolbox developed by the HotQCD collaboration Clarke:2023sfy . We thank Luis Altenkort for providing his analysis code luhuhis . The authors acknowledge support from the Deutsche Forschungsgemeinschaft (DFG) through the CRC-TR 211 "Strong-interaction matter under extreme conditions" (Project No. 315477589 - TRR 211). J.H.W.’s research is also funded by the DFG under project number 417533893/GRK2575 "Rethinking Quantum Field Theory". J.H.W. acknowledges the support by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). R.L. was supported by the Ministry of Culture and Science of the State of Northrhine Westphalia (MKW NRW) under the funding code NW21-024-A (NRW-FAIR). This research used computing time provided by the ALCC and INCITE programs at the Oak Ridge Leadership Computing Facility, a DOE Office of Science User Facility supported under contract no. DE-AC05-00OR22725; the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility at Lawrence Berkeley National Laboratory under Contract No. DE-AC02-05CH11231; PRACE awards on JUWELS at GCS@FZJ, Germany; Marconi100 at CINECA, Italy; and the LUMI-G supercomputer at the Finnish IT Center for Science (CSC). Computations were also performed on facilities of the USQCD Collaboration funded by the DOE Office of Science, in particular on the 21-g cluster at the Thomas Jefferson National Accelerator Facility (JLAB). In addition, parts of this work were performed on the GPU cluster of the University of Bielefeld, supported by HPC.NRW. Data Availability Statement: The raw and derived data of this study are openly available in bollweg2025data . The gauge field configurations are available on request and will be openly available on the International Lattice Data Grid (ILDG) at a later stage. Code Availability Statement: The lattice generation and GF analysis presented in this work were performed using SIMULATeQCD HotQCD:2023ghu and the MILC code. The analysis code is openly available in AnalysisToolbox ; luhuhis ; bollweg2025data .